# Homogeneous Tokenizer Matters: Homogeneous Visual Tokenizer for Remote Sensing Image Understanding

### v1

@@ B. Subpatch-level methods
Subpatch-level methods argue that the implicit assumption in Patch Embed, where all patches in an image are treated equally, is unreasonable. These methods proposed that the patches corresponding to foreground regions in the image should be refined, meaning that more patches should be used to describe important areas of the image.
W. Chen et al. utilized class activation maps (CAMs) to identify salient regions in images and proposed using smaller patches for these significant areas to increase the overall number of patches[34]. Similarly, T. Ronen et al. employed GradCAM to locate salient regions and guided the generation of fine-grained patches using quadtree structures[35]. M. Chen et al. introduced CF-ViT[36] and used the confidence levels from a coarse-grained model to determine whether patches should undergo further processing. Y. Wang et al. also leveraged results from a coarse-grained model, continually densifying patches across the image until high-confidence recognition outcomes were obtained[37]. X. Yue et al. proposed PS-ViT[38], which employs progressive sampling to shift token positions toward the main regions of interest in the image, enabling a more detailed description of foreground elements through additional tokens.
@@ C. Superpatch-level methods
Superpatch-level methods argue that the use of fixed-size patches restricts the model to capturing information at a single scale and propose the construction of multiscale patches to enhance the model’s robustness across different scales.
C.-F. Chen et al. introduced CrossViT[39], which uses two patches of different sizes to split an image, enabling the extraction of multiscale features from the same image. L. Beyer et al. proposed FlexiViT[40], which enhances the model’s robustness to different patch sizes by randomly interpolating parameters in the Patch Embed layer. S. H. Lee et al. presented shifted patch tokenization, which applies multiangle shifting and aggregation of patches to address the lack of local inductive bias in Patch Embed[41].
2.2 Object-oriented Visual Tokenizer
While the above methods have made significant progress in improving Patch Embed, patch-based approaches still cannot fundamentally overcome the limitations of rectangular patches. Currently, some research has begun to explore moving away from patches and instead reconstructing visual tokenizers based on the objects in the images.
B. Wu et al. proposed VT[42], which employs a simple convolutional operation to assign each pixel in an image to one of several semantic groups, mapping them to visual tokens. However, VT lacks global semantic awareness of the image, potentially leading to an overreliance on local pixel characteristics such as color and texture. T. Yang et al. introduced visual concept tokenization (VCT)[43], which utilizes concept token reconstruction in a cleverly designed pretraining task to enable individual control over specific features by each visual token on simulated images. However, VCT requires pretraining and features a complex model structure. S. Qian et al. developed MoTo[23], which perceives the semantic consistency association between tokens but does not represent them as homogeneous visual tokens. J. Mei et al. proposed SPFormer[44], which divides images into irregular, semantically homogenous regions. However, the method relies on superpixel algorithms and lacks plug-and-play functionality.
To fundamentally address the limitations of the aforementioned approaches, we revisited tokenization in natural language processing and proposed two fundamental properties that a visual tokenizer should possess: homogeneity and adaptability. Building upon this premise, we introduced the HOmogeneous visual tOKenizer (HOOK), which demonstrates distinct advantages in both performance and efficiency over baselines, including Patch Embed. HOOK has the potential to replace patch-based methods as a new foundational visual tokenizer.
@@ 3 Theoretical Analysis
In this section, we conducted a theoretical analysis of visual tokens. In Section 3.1, we emphasized the importance and necessity of homogeneity in visual tokenizers by examining the essence of tokenization. In Section 3.2, we defined the confusion matrix for tokens and objects under strict definitions and constraints. Section 3.3 proposed two general routings for constructing homogeneous visual tokens based on the confusion matrix that was previously established.
@@ 3.1 Why do we need a homogeneous visual tokenizer?
The concept of the tokenizer originates from the field of natural language processing (NLP). Machines are unable to understand unstructured data, and the first step in enabling machines to comprehend language is to structurize the unstructured language documents. The NL tokenizer is one of the most popular methods for achieving this process by breaking down language into smaller basic elements known as tokens.
A. NL tokenizer
In the NL tokenizer, the definition of a token is not fixed. Following the order from larger tokens to smaller tokens, a natural language tokenizer can be classified into a sentence tokenizer, word tokenizer, subword tokenizer, and character tokenizer, among others. The size of the token involves a balance between efficiency and generalization:
@@ A. NL tokenizer
(2) Smaller tokens have stronger generalizability but are less efficient. For instance, in a character tokenizer, the vocabulary consists of only 26 letters, with necessary symbols. Any new word can be composed of these characters. In theory, a character tokenizer can generalize to any text. However, because individual letters lack semantic meaning, machines struggle to learn words, sentences, and higher-level semantic information. Moreover, several tokens may be required to represent just one word.
Therefore, NL tokenizers need to define suitable basic elements of language to balance the conflicting goals of efficiency and generalization. The most popular NL tokenizers, such as Word2Vec[13], BPE[45], WordPiece[46], Unigram[16], and others[14, 15] belong to the category of word tokenizers or subword tokenizers, which use words or subwords as the basic elements of language. Words and subwords represent the smallest semantically independent elements in language, effectively striking a balance between efficiency and generalization.
B. Visual tokenizer
Research on visual tokenization has been driven by transformer-based vision models, which are aimed at identifying the basic elements of an image and splitting the image into a token sequence to meet the input requirements of transformer models. Like NL tokenizers, visual tokenizers also face the challenge of balancing efficiency and generalizability. Consider two extreme scenarios:
(1) Tokens that represent individual pixels. This approach offers the best generalizability because it can generalize to any image within the same color space. However, this method has the lowest efficiency: it requires 50,176 tokens to represent just one 224 $\times$ 224 image.
@@ B. Visual tokenizer
In balancing these aspects, visual tokenizers are aimed at finding a suitable granularity of tokens that balances efficiency and generalization for effectively processing images within transformer-based models.
An intuitive approach is to find an intermediate value between the two extreme scenarios mentioned above, similar to words and subwords in NL tokenizers. Notably, the most popular patch-based methods align with this approach. Using Patch Embed as an example, it uses a size of 16 $\times$ 16 as an empirical middle ground to balance efficiency and generalizability. This is a fundamental reason why patch-based methods are prevalent in visual tokenization.
@@
C. Semantically Independent Region
As shown in Figure 1-d, based on the observations and theoretical analysis mentioned above, we have defined the semantically independent region (SIR) as a replacement for the fixed-size rectangular patch as the basic element of an image. SIRs refer to regions in the image that are semantically independent of other regions outside themselves. Once semantically independent regions are defined, they are internally homogeneous in semantics compared to inter-regions. This is why we believe that an ideal visual tokenizer should possess homogeneity.
In addition, in the definition of SIRs, ”independence” is a dynamic concept. For example, an entire aircraft is an independent region in relation to the ground, wings are independent regions in relation to the aircraft, and flaps are independent regions in relation to the wings. Therefore, when addressing a specific remote sensing image, the image size and the granularity of the task determine the ideal granularity of independence, while the number of tokens determines the actual granularity of independence. Additionally, a visual tokenizer should be image-agnostic and task-agnostic. Therefore, the number of tokens should be adjusted to accommodate images of any size and tasks of any granularity. We refer to this property as the adaptivity of the visual tokenizer.
@@ 3.2 Confusion matrix for tokens and objects
Figure 2: Confusion matrix of tokens and objects. In a simplified scenario and under strict definitions, the confusion matrix of tokens and objects reveals two general routings for constructing homogeneous visual tokens: (1) splitting and merging and (2) merging and splitting.
Figure 1-c illustrates the confusion matrix between tokens and objects. In this section, we define this confusion matrix under more rigorous definitions and constraints.
First, we simplify the scenario: as shown in the left image of each set of images in Figure 2, we assume that a remote sensing image contains only two objects, and we are not concerned with the background. The tokens corresponding to the background regions are referred to as “background tokens”.
Second, we define the following concept:
@@
“Region of the token” refers to the pixel area on the original image that contains the semantic information encapsulated by a specific token.
For ease of articulation and accuracy, we further define two types of relationships between objects and regions of the tokens:
@@
“Cover” refers to the complete inclusion of an object in the region of the token on the image.
@@
“Overlap” refers to the presence of an intersection between the region of the token and an object.
From the above definitions, it is evident that if the region of the token covers an object, they must overlap, and the reverse is not true.
Based on the above definitions, we can provide precise definitions for the four scenarios mentioned in Figure 1-c and Figure 2:
@@
”Same Object Same Token” means that, excluding background tokens, any region of the token covers and only covers one object, and each region of the token is distinct from other regions.
@@
”Same Object Multiple Tokens” means that no region of the token covers an object, and at most, the region overlaps one object.
@@
”Same Token Multiple Objects” means that, excluding background tokens, each region of the token covers at least two objects.
@@
”Multiple Tokens Multiple Objects” means that there is a region of the token that overlaps with two objects but covers at most one object.
As shown in Figure 2, the above definition completely describes the relationship between tokens and objects.
3.3 General routing to homogeneous visual tokens
In Section 1, we show that “Same Object Multiple Tokens” makes it difficult for the model to learn complete object features, that “Same Token Multiple Object” makes it difficult for the model to learn the relationships between objects, and “Multiple Tokens Multiple Objects” inherits the drawbacks of the above two cases. The ideal scenario between tokens and objects should be “Same Object Same Token”, and we refer to this token as a homogeneous visual token. In other words, in the four squares of Figure 2, only square 1 is ideal. Because the default visual tokenizer, Patch Embed, belongs to square 4, constructing the ideal visual token can be abstracted as a routing selection problem from square 4 to square 1.
According to the confusion matrix in Figure 2, without considering diagonal routings, each step has only two choices: ”move left” and ”move up.” ”Move left” refers to keeping the same number of objects but reducing the number of tokens from multiple to one. For example, going from ② to ① signifies that multiple tokens corresponding to one object are reduced to one token corresponding to one object, with the basic operation for this step being to “merge” tokens. “Move up” signifies keeping the same number of tokens but reducing the number of objects from multiple to one. For example, going from ③ to ① signifies that one token corresponding to multiple objects is reduced to one token corresponding to one object, with the basic operation for this step being to “split” tokens.
Thus, we have obtained two general routings for constructing homogeneous visual tokens:
@@ 4 Method
Section 3.3 outlined two general routings for constructing homogeneous visual tokens. Based on the ”Splitting and Merging” routing, we designed a simple HOmogeneous visual tOKener: HOOK. In Section 4.1, we provide an overview of the structure of HOOK. Section 4.2 and Section 4.3 detail two important modules in HOOK: the object perceiver module and the object vectorization module. In Section 4.4, we discussed how HOOK, as a visual tokenizer, adapts to images of any size and tasks of any granularity.
@@ 4.1 Overview
Figure 3: The architecture of HOOK, which mainly consists of two modules: the object perception module (OPM), which is responsible for perceiving semantically independent regions, and the object vectorization module (OVM), which is responsible for vectorizing semantically independent regions into tokens.
The overall architecture of HOOK is shown in Figure 3. HOOK can be viewed as a function $T$ that maps the image $I$ to $N$ D-dimensional vectors, with each vector representing a token, as follows:
@@
where $t\in R^{N\times D}$ and $I\in R^{H\times W\times C}$.
HOOK primarily consists of two modules: the object perception module (OPM) and the object vectorization module (OVM).
The OPM is aimed at perceiving semantically independent regions within the image. Specifically, the OPM initially splits the image into several 4 $\times$ 4 pixel-sized seeds using convolutional blocks, followed by stacked local and global self-attention layers to expand the seeds into semantically independent regions (as detailed in Section 4.2). The process above can be formalized as follows:
@@
where $P$ represents the object perception module and $f\in R^{\frac{HW}{16}\times d}$ represents the seeds after passing through the self-attention layers, where $d$ denotes the dimension of the seeds.
The OVM is aimed at vectorizing semantically independent regions into tokens while achieving arbitrary adjustment of the token quantity. Specifically, we define $N$ learnable vectors $q$ as queries and utilize the cross-attention mechanism to merge seeds belonging to the same semantically independent region into homogeneous tokens. Here, $N$ acts as a variable hyperparameter to enable arbitrary adjustment of the token quantity (as detailed in Section 4.3). The process above can be formalized as follows:
@@
where $q\in R^{N\times D}$ and $V$ represents the object vectorization module.
In conclusion, the entire tokenizer is formalized as:

### v2

@@ B. Subpatch-level methods
Subpatch-level methods are based on the argument that the implicit assumption in Patch Embed, namely, that all patches in an image should be treated equally, is unreasonable. In these methods, it is assumed that the patches corresponding to foreground regions in the image should be refined, meaning that more patches should be used to describe important areas of the image.
W. Chen et al. utilized class activation maps (CAMs) to identify salient regions in images and proposed the use of smaller patches for these significant areas to increase the overall number of patches[36]. Similarly, T. Ronen et al. employed GradCAM to locate salient regions and guide the generation of fine-grained patches via quadtree structures[37]. M. Chen et al. introduced CF-ViT[38] and used confidence levels from a coarse-grained model to determine whether patches should undergo further processing. Y. Wang et al. also leveraged results from a coarse-grained model, continually densifying patches across the image until high-confidence recognition outcomes were obtained[39]. X. Yue et al. proposed PS-ViT[40], which employs progressive sampling to shift token positions towards the main regions of interest in the image, enabling a more detailed description of foreground elements through additional tokens.
@@ C. Superpatch-level methods
Superpatch-level methods are based on the argument that the use of fixed-size patches restricts the model to capturing information at a single scale. These methods construct multiscale patches to increase the model’s robustness across different scales.
C.-F. Chen et al. introduced CrossViT[41], which uses two patches of different sizes to split an image, enabling the extraction of multiscale features from the image. L. Beyer et al. proposed FlexiViT[42], which enhances the model’s robustness to different patch sizes by randomly interpolating parameters in the Patch Embed layer. S. H. Lee et al. presented shifted patch tokenization, which applies multiangle shifting and aggregation of patches to address the lack of local inductive bias in Patch Embed[43].
2.2 Object-Oriented Visual Tokenizers
While the above methods have made significant progress in improving Patch Embed, patch-based approaches still cannot fundamentally overcome the limitations of rectangular patches. Researchers have begun to explore moving away from patches and instead reconstructing visual tokenizers on the basis of the objects in the images.
B. Wu et al. proposed VT[44], which employs a simple convolution operation to assign each pixel in an image to one of several semantic groups, mapping them to visual tokens. However, VT lacks global semantic awareness of the image, potentially leading to an overreliance on local pixel characteristics such as colour and texture. T. Yang et al. introduced visual concept tokenization (VCT)[45], which uses concept token reconstruction in a cleverly designed pretraining task to enable individual control over specific features by each visual token on simulated images. However, VCT requires pretraining and features a complex model structure. S. Qian et al. developed MoTo[25], which perceives the semantic consistency association between tokens but does not represent them as homogeneous visual tokens. J. Mei et al. proposed SPFormer[46], which divides images into irregular, semantically homogenous regions. However, the method relies on superpixel algorithms and lacks plug-and-play functionality.
To address the limitations of the aforementioned approaches fundamentally, we revisit tokenization in natural language processing and propose two fundamental properties that a visual tokenizer should possess: homogeneity and adaptability. Building upon this premise, we introduce the HOmogeneous visual tOKenizer (HOOK), which has distinct advantages in terms of both performance and efficiency over baselines, including Patch Embed. HOOK has the potential to replace patch-based methods as a new foundational visual tokenizer.
@@ 3 Theoretical Analysis
In this section, we conduct a theoretical analysis of visual tokens. In Section 3.1, we emphasize the importance and necessity of homogeneity in visual tokenizers by examining the essence of tokenization. In Section 3.2, we define a confusion matrix for tokens and objects under strict definitions and constraints. In Section 3.3, we propose two general routes for constructing homogeneous visual tokens on the basis of the confusion matrix that was previously established.
@@ 3.1 Why do we need a homogeneous visual tokenizer?
The concept of a tokenizer originates from the field of natural language processing (NLP). Machines are unable to understand unstructured data, and the first step in enabling machines to comprehend language is to structurize unstructured language documents. NL tokenizers are among the most popular methods for achieving this process by breaking down language into smaller basic elements known as tokens.
A. NL tokenizers
In an NL tokenizer, the definition of a token is not fixed. Following the order from larger to smaller tokens, natural language tokenizers can be classified into sentence tokenizers, word tokenizers, subword tokenizers, and character tokenizers, among others. The size of the tokens involves a balance between efficiency and generalization:
@@ A. NL tokenizers
(2) Smaller tokens have stronger generalizability but are less efficient. For example, in a character tokenizer, the vocabulary consists of only 26 letters, with necessary symbols. Any new word can be composed of these characters. In theory, a character tokenizer can be generalized to any text. However, because individual letters lack semantic meaning, machines struggle to learn words, sentences, and higher-level semantic information. Moreover, several tokens may be required to represent just one word.
Therefore, NL tokenizers need to define suitable basic elements of language to balance the conflicting goals of efficiency and generalization. The most popular NL tokenizers, such as Word2Vec[13], Byte-Pair Encoding (BPE)[47], WordPiece[48], Unigram[16], and others[14, 15], belong to the category of word tokenizers or subword tokenizers, which use words or subwords as the basic elements of language. Words and subwords represent the smallest semantically independent elements in language, effectively striking a balance between efficiency and generalization.
B. Visual tokenizers
Research on visual tokenization has been driven by transformer-based visual models, which are aimed at identifying the basic elements of an image and splitting the image into a token sequence to meet the input requirements of transformer models. Like NL tokenizers, visual tokenizers also face the challenge of balancing efficiency and generalizability. We consider two extreme scenarios:
(1) Tokens that represent individual pixels. This approach offers the best generalizability because it can be generalized to any image within the same colour space. However, this method has the lowest efficiency: it requires 50,176 tokens to represent just one 224 $\times$ 224 image.
@@ B. Visual tokenizers
In balancing these aspects, visual tokenizers are aimed at finding a suitable granularity of tokens that balances efficiency and generalizability for effectively processing images within transformer-based models.
An intuitive approach is to find an intermediate value between the two extreme scenarios mentioned above, similar to words and subwords in NL tokenizers. Notably, the most popular patch-based methods align with this approach. For example, Patch Embed uses patches of size 16 $\times$ 16 as an empirical middle ground to balance efficiency and generalizability. This is a fundamental reason why patch-based methods are prevalent in visual tokenization.
@@
C. Semantically Independent Regions
As shown in Figure 1-d, on the basis of the observations and theoretical analysis discussed above, we define the semantically independent region (SIR) as a replacement for the fixed-size rectangular patch as the basic element of an image.
Definition 1
Semantically independent region (SIR) refers to a distinct and self-contained region within an image that has a specific semantic meaning and is distinguishable from its surroundings.
An SIR typically includes a single visual object, such as an airplane, a car, or a building, and it visually stands out from the surrounding environment and semantically represents a complete conceptual entity. Once semantically independent regions are defined, they are internally homogeneous in terms of semantics compared with interregions. This is why we believe that an ideal visual tokenizer should possess homogeneity.
In addition, in the definition of SIRs, ”independence” is a dynamic concept. For example, an entire aircraft is an independent region in relation to the ground, wings are independent regions in relation to the aircraft, and flaps are independent regions in relation to the wings. Therefore, when addressing a specific remote sensing image, the image size and granularity of the task determine the ideal granularity of independence, whereas the number of tokens determines the actual granularity of independence. Additionally, an ideal visual tokenizer should be image-agnostic and task-agnostic. Therefore, the number of tokens should be adjusted to accommodate images of any size and tasks of any granularity. We refer to this property as the adaptivity of the visual tokenizer.
@@ 3.2 Confusion matrix for tokens and objects
Figure 2: Confusion matrix of tokens and objects. In a simplified scenario and under strict definitions, the confusion matrix of tokens and objects reveals two general routes for constructing homogeneous visual tokens: (1) splitting and merging and (2) merging and splitting.
Figure 1-c illustrates the confusion matrix between tokens and objects in four real remote sensing images. To analyse the relationships among the four scenarios in the confusion matrix more clearly, we abstract and simplify the real scenes.
As shown in Figure 2, we assume that each remote sensing image contains only two objects and that we are not concerned with the background. The tokens corresponding to the background regions are referred to as ”background tokens”. In fact, real remote sensing images containing multiple complex objects can be extrapolated from the simplified scenarios described above.
For rigour of presentation, we define the following concept:
@@
”Region of the token” refers to the pixel region on the original image that contains the semantic information encapsulated by a specific token.
For ease of presentation and accuracy, we further define two types of relationships between the objects and regions of the tokens:
@@
”Cover” refers to the complete inclusion of an object in the region of the token on the image.
@@
”Overlap” refers to the presence of an intersection between the region of the token and an object.
From the above definitions, if the region of the token covers an object, they must overlap, whereas the reverse is not true.
On the basis of the above definitions, we can provide precise definitions for the four scenarios mentioned in Figure 1-c and Figure 2:
@@
”Same object same token” means that, excluding background tokens, any region of the token covers exactly one object, and each region of the token is distinct from other regions.
@@
”Same object multiple tokens” means that no region of a token covers an object, and at most, each region overlaps with one object.
@@
”Same token multiple objects” means that, excluding background tokens, each region of a token covers at least two objects.
@@
”Multiple tokens multiple objects” means that there is a region of each token that overlaps with two objects but covers at most one object.
As shown in Figure 2, the above definitions completely describe the relationships between tokens and objects.
3.3 General route to homogeneous visual tokens
In Section 1, we show that ”same object multiple tokens” makes it difficult for the model to learn complete object features, that ”same token multiple objects” makes it difficult for the model to learn the relationships between objects, and that ”multiple tokens multiple objects” inherits the drawbacks of the above two cases. The ideal scenario between tokens and objects should be ”same object same token”, and we refer to this token as a homogeneous visual token. In other words, in the four squares of Figure 2, only square 1 is ideal. Because the default visual tokenizer, Patch Embed, corresponds to square 4, constructing the ideal visual token can be abstracted as a route selection problem from square 4 to square 1.
According to the confusion matrix in Figure 2, without considering diagonal routes, each step has only two choices: ”move left” and ”move up.” ”Move left” refers to keeping the same number of objects but reducing the number of tokens from multiple to one. For example, going from ② to ① signifies that multiple tokens corresponding to one object are reduced to one token corresponding to one object, with the basic operation for this step being to ”merge” tokens. ”Move up” signifies keeping the same number of tokens but reducing the number of objects from multiple to one. For example, going from ③ to ① signifies that one token corresponding to multiple objects is reduced to one token corresponding to one object, with the basic operation for this step being to ”split” tokens.
Thus, we obtain two general routes for constructing homogeneous visual tokens:
@@ 4 Method
Section 3.3 outlines two general routes for constructing homogeneous visual tokens. On the basis of the ”splitting and merging” route, we design a simple HOmogeneous visual tOKenizer: HOOK. In Section 4.1, we provide an overview of the structure of HOOK. Section 4.2 and Section 4.3 detail two important modules in HOOK: the object perception module and the object vectorization module. In Section 4.4, we discuss how HOOK adapts to images of any size and tasks of any granularity.
@@ 4.1 Overview
Figure 3: The architecture of HOOK consists of two modules: the object perception module (OPM), which is responsible for perceiving semantically independent regions, and the object vectorization module (OVM), which is responsible for vectorizing semantically independent regions into tokens.
The overall architecture of HOOK is shown in Figure 3. HOOK can be viewed as a function $T$ that maps an image $I$ to $N$ D-dimensional vectors, with each vector representing a token, as follows:
@@
where $t\in\mathbb{R}^{N\times D}$ and $I\in\mathbb{R}^{H\times W\times C}$.
HOOK consists of two modules: the object perception module (OPM) and the object vectorization module (OVM).
The OPM is aimed at perceiving semantically independent regions within the image. Specifically, the OPM initially splits the image into several 4 $\times$ 4 pixel-sized seeds via convolutional blocks, which are followed by stacked local and global self-attention layers to expand the seeds into semantically independent regions (as detailed in Section 4.2). The process above can be formalized as follows:
@@
where $P$ represents the object perception module and where $f\in\mathbb{R}^{\frac{HW}{16}\times d}$ represents the seeds after passing through the self-attention layers, where $d$ denotes the dimension of the seeds.
The OVM is aimed at vectorizing semantically independent regions into tokens while achieving arbitrary adjustment of the token quantity. Specifically, we define $N$ learnable vectors $q$ as queries and utilize the cross-attention mechanism to merge seeds belonging to the same semantically independent region into homogeneous tokens. Here, $N$ acts as a variable hyperparameter to enable arbitrary adjustment of the token quantity (as detailed in Section 4.3). This process can be formalized as follows:
@@
where $q\in\mathbb{R}^{N\times D}$ and where $V$ represents the object vectorization module.
In conclusion, the entire tokenizer is formalized as